# A Matter of Annotation: An Empirical Study on In Situ and Self-Recall Activity Annotations from Wearable Sensors

### v2

@@ Abstract
Research into the detection of human activities from wearable sensors is a highly active field, benefiting numerous applications, from ambulatory monitoring of healthcare patients via fitness coaching to streamlining manual work processes. We present an empirical study that compares 4 different commonly used annotation methods utilized in user studies that focus on in-the-wild data. These methods can be grouped into user-driven, in situ annotations - which are performed before or during the activity is recorded - and recall methods - where participants annotate their data in hindsight at the end of the day. Our study illustrates that different labeling methodologies directly impact the annotations’ quality, as well as the capabilities of a deep learning classifier trained with the data. We noticed that in situ methods produce less but more precise labels than recall methods. Furthermore, we combined an activity diary with a visualization tool that enables the participant to inspect and label their activity data. Due to the introduction of such a tool were able to decrease missing annotations and increase the annotation consistency, and therefore the F1-Score of the deep learning model by up to 8% (ranging between 82.1 and 90.4 % F1-Score). Furthermore, we discuss the advantages and disadvantages of the methods compared in our study, the biases they could introduce, and the consequences of their usage on human activity recognition studies as well as possible solutions.
@@
Sensor-based activity recognition is one of the research fields of Pervasive Computing developed with enormous speed and success by industry and science and influencing medicine, sports, industry, and therefore the daily lives of many people. However, current available smart devices are mostly capable of detecting periodic activities like simple locomotions. To recognize more complex activities a multimodal sensor input, such as [1], and more complex recognition models are needed. Many of the published datasets are made in controlled laboratory environments. Such data does not have the same characteristics and patterns as data recorded in-the-wild. Data that belongs to similar classes but is recorded in an uncontrolled versus controlled environment can differ significantly since it contains more contextual information [2]. Furthermore, study participants tend to control their movements more while being monitored [3]. The recording of long-term and real-world data is a tedious, time-consuming, and therefore a non-trivial task. Researchers have various motivations to record such datasets but the technical hurdles are still high and problems during the annotation process occur regularly. With regards to Human Activity Recognition (HAR), recording a long-term dataset always presents the researcher with the problem of developing a methodology that, on one hand, allows precise labels and, on the other hand, does not unnecessarily burden or disturb the study participants. Relying only on self-recall methods, like writing an activity diary, e.g. [4], can result in imprecise time indications that do not necessarily correspond to the actual periods of an activity. Such incorrectly or noisy labeled data [5] leads to a trained model that is less capable of detecting activities reliably, due to unwanted temporal dependencies learned by wrongly annotated patterns [6].
We can see an emerging spotlight on real-world and long-term activity recognition and think that it will be one of the main research challenges that need to be put more in focus to overcome current limitations and be capable of recognizing complex day-to-day activities.
Such datasets, rely heavily on self-recall methods or using additional apps to track movements and set labels either automatically [7] or with the manual selection of a label [8]. Due to these hurdles, many researchers prefer to work with datasets from controlled over uncontrolled environments. As a consequence, only a very limited number of in-the-wild datasets have been published until now.
@@
Furthermore, we used a Shallow-DeepConv(LSTM) architecture, see Bock et al. [10] and Ordoñez et al. [11], and trained models with a leave-one-day-out cross-validation method of 6 previously selected subjects and each annotation method.
@@
1 Materials and Methods
@@
Our study is conducted with 11 participants, from which 10 are male and 1 is female. The participants are between 25 and 45 years old. Out of 11 participants, 6 are researchers in the field of signal processing and are used to read and work with sensor data. Participants were selected among acquaintances and colleagues. The study was conducted over 2 weeks while participants wore an open-source smartwatch on their wrist of choice. During the two-week study, the participants were instructed to use 4 different labeling methods in parallel, see Figure 1. In the first week they were asked to use the in situ button, in situ app, and pure self-recall methods. At the beginning of the 2nd week, we expanded the number of annotation methods with the time-series recall. This annotation method combines the activity diary with a graphical visualization of the participants’ daily data.
The Bangle.js smartwatch has 3 mechanical buttons on the right side of the case. These buttons are programmed to record the number of consecutive button presses per minute. The total number of button presses is stored with the given timestamp and can therefore be used to mark the beginning and end of an activity in the time-series.
@@
The time-series recall method can be seen as a combination of an activity diary and a graphical representation of the raw sensor data. For visualization and labeling purposes, we provided the participants with an adapted version of the MAD-GUI. The GUI was published by Ollenschläger et al. [9] in 2022 and is a generic open-source Python package. Therefore, it can be integrated into one’s project. Our adaptions to the package are available for download from a GitHub repository333https://github.com/ahoelzemann/mad-gui-adaptions/. It contains changes to the data loader, the definition of available labels, and color settings for displaying the 3D raw data.
Annotation Guidelines: The participants were briefed to note daily returning activities (sports or activities of daily living) that are performed longer than 10 minutes roughly 2 hours after the recording stops. The name of the activity was chosen individually by the participant.
Participants decided individually at what time of the day the recording would start on the next day.
Each of these annotation methods represents a layer of annotation that is used for the visual, statistical, and deep learning evaluation. Figure 1 illustrates the overall concept.
@@ 2 Statistical Analysis
The labels were statistically analyzed based on their consistency using the Cohen score as well as the number of missing annotations across all methods. The Cohen score describes the agreement between two annotation methods, which is defined as follows (see [12] and [13]). Where is the observed agreement ratio and is the expected agreement if both annotators assign labels randomly. The score shows how uniform two different annotators labeled the same data. For calculation purposes, an implementation provided by Scikit-Learn [14], was used. Furthermore, missing annotations across methods are measured as the percentage of missing or incomplete annotations. The annotations of all methods were first compared with each other and matched based on the given time indications. Annotations that could not be assigned or were missing were marked accordingly and are the base for calculating this indicator, Figure 4 visualizes this. We used a similar representation as [15] to visualize the matches among labeling methods. In this study, the authors compared genome annotations labeled by different annotators with regard to their error scores between different annotators.
@@ 3 Effects on Deep Learning Performance
The implementation of [16] incorporates BatchNormalization layers after each Convolutional layer, as well as MaxPooling for the transition between the final convolutional block and the LSTM layer, and a Dropout layer before classification. Each Convolutional layer employs a ReLU activation function. The inclusion of the BatchNormalization layers serves to accelerate training and mitigate the detrimental effects of internal covariate shift, as discussed further in [21]. Before training the neural network we apply two preprocessing steps to the data. These are an upsampling to a constant 25 Hz due to small deviations in the device sampling rate as well as a rescaling of the accelerometer data between -1 and 1.
Leave-One-Day-Out Cross-Validation: Figure 3 shows the train- and test-setting for the deep learning model. For every study participant and every week, a personalized model has been trained and tested.
@@
We adapted the leave-one-subject-out strategy to fit our needs. Instead of using one participant for testing, we used one day of the week for testing and trained on every other day. The dataset consists of a major void-class and a small number of samples per activity class and participant. To counteract an above-average large void-class, we trained our model with balanced class weights. Not limiting the participants in their choice of daily performed activities as well as well as not specifying predefined activity labels resulted in very unique sets of activities per study participant. Due to these circumstances, we cannot expect that it is possible to train a model that is capable of generalizing across participants and days. Every participant comes with their specific patterns to perform an activity. Furthermore, due to the in-the-wild recording setup, the intra-class differences [22] for comparatively simple activities, walking or running can be significant. The impact of different labeling methods is therefore expected to be more present, and hence more visible, in a personalized than it would in a generalized model.
Postprocessing & Classification: We classified the data based on a sliding window with a length of 2 seconds (50 samples). However, we are looking for longer periods of reoccurring activities and decided therefore to apply a jumping window of 5 minutes that applies a majority vote to the given time period. The activity with the most instances in those 5 minutes is set for the whole window.
@@ 4 Results
Our participants were asked to annotate activities carried out in their daily lives performed for more than 10 min. We didn’t limit the participants to a predefined set of classes. They decided independently which labels they would like to set for certain periods. After normalizing the labels, e.g. changing “going for a walk” to “walking” etc., the set of labels as given by the participants contains the following 23 (22 + void) different labels: laying, sitting, walking, running, cycling, bus_driving, car_driving, vacuum_cleaning, laundry, cooking, eating, shopping, showering, yoga, sport, playing_games, desk_work, guitar_playing, gardening, table_tennis, badminton, horse_riding. Every sample that wasn’t specifically labeled is classified as void.

### v3

@@ Abstract
Research into the detection of human activities from wearable sensors is a highly active field, benefiting numerous applications, from ambulatory monitoring of healthcare patients via fitness coaching to streamlining manual work processes.
We present an empirical study that evaluates and contrasts four commonly employed annotation methods in user studies focused on in-the-wild data collection. For both the user-driven, in situ annotations, where participants annotate their activities during the actual recording process, and the recall methods, where participants retrospectively annotate their data at the end of each day, the participants had the flexibility to select their own set of activity classes and corresponding labels.
Our study illustrates that different labeling methodologies directly impact the annotations’ quality, as well as the capabilities of a deep learning classifier trained with the data. We noticed that in situ methods produce less but more precise labels than recall methods. Furthermore, we combined an activity diary with a visualization tool that enables the participant to inspect and label their activity data. Due to the introduction of such a tool were able to decrease missing annotations and increase the annotation consistency, and therefore the F1-Score of the deep learning model by up to 8% (ranging between 82.1 and 90.4 % F1-Score). Furthermore, we discuss the advantages and disadvantages of the methods compared in our study, the biases they could introduce, and the consequences of their usage on human activity recognition studies as well as possible solutions.
@@
sectionIntroduction
Sensor-based activity recognition is one of the research fields of Pervasive Computing developed with enormous speed and success by industry and science and influencing medicine, sports, industry, and therefore the daily lives of many people. However, current available smart devices are mostly capable of detecting periodic activities like simple locomotions. In order to recognize more complex activities a multimodal sensor input, such as [1], and more complex recognition models are needed. Many of the published datasets are made in controlled laboratory environments. Such data does not have the same characteristics and patterns as data recorded in-the-wild. Data that belongs to similar classes but is recorded in an uncontrolled versus controlled environment can differ significantly since it contains more contextual information [2] . Furthermore, study participants tend to control their movements more while being monitored [3]. The recording of long-term and real-world data is a tedious, time-consuming, and therefore a non-trivial task. Researchers have various motivations to record such datasets but the technical hurdles are still high and problems during the annotation process occur regularly.
In Human Activity Recognition research, capturing long-term datasets presents a challenge: balancing precise labeling with minimal participant burden. Sole reliance on self-recall methods, like activity diaries e.g., [4], often leads to imprecise time indications that may not accurately reflect actual activity duration. Such incorrectly or noisy labeled data later on leads to a trained model that is less capable of detecting activities reliably [5], due to unwanted temporal dependencies learned by wrongly annotated patterns [6].
The field of HAR is witnessing a growing emphasis on real-world and long-term activity recognition. This focus stems from the need to address current limitations and achieve reliable recognition of complex daily activities. Existing long-term datasets often rely heavily on self-recall methods or additional tracking apps. These apps can set labels either automatically [7] or require manual selection [8]. However, such approaches present challenges, leading many researchers to favor controlled environments for data collection. As a consequence, the number of publicly available ”in-the-wild” datasets remains limited.
@@
The participants in this study were given the freedom to self-report their activity classes based on the diverse range of pursuits encompassed within their daily lives. Consequently, the resulting dataset exhibits a heterogeneous composition, comprising both commonplace, routine activities such as walking, driving, and eating, as well as more specialized and niche activities like badminton, yoga, horse_riding, and gardening.
Furthermore, we used a Shallow-DeepConv(LSTM) architecture, see [10] and [11], and trained models with a Leave-One-Day-Out cross-validation method of 6 previously selected subjects and each annotation method.
@@
1 Related work
A very limited number of datasets are currently publicly available which were recorded in the wild, e.g. [12, 13, 14, 15, 16]. [16] and [15] were captured in naturalistic settings, but the participants were equipped with multiple sensors on various body locations and were filmed by a third party during the exercises. Such visible equipment and the presence of an observer could potentially introduce a behavior bias [17] in the data, as it may alter participants’ movement patterns due to the constant reminder that they are participating in a study [3]. Furthermore, multimodal datasets recorded with multiple body-worn sensors, rather than a single Inertial Measurement Unit (IMU), have faced the challenge of proper inter-sensor synchronization [18]. A comprehensive dataset encompassing a diverse range of classes, accurate annotations, and recorded by a single device that is nearly unnoticeable to the participant (and therefore unlikely to influence their behavior or movement patterns), is not yet publicly available due to the aforementioned obstacles.
According to [19] and later [8], we distinguish between 6 or 7, respectively, different methods and 2 environments (online/offline) of labeling data, the methods are (1) Indirect Observation, (2) Self-Recall, (3) Experience Sampling, (4) Video/Audio Recordings, (5) Time Diary, (6) Human Observer, (7) Prompted Labeling. [20] uses 4 different categories to classify data labeling approaches, these are (1) temporal (when) - is the label conducted during or after the activity, (2) annotator (who) - is the label given by the individual itself or by an observer, (3) scenario (where) - is the activity labeled in a controlled (e.g laboratory) or uncontrolled (in-the-wild) environment, and (4) annotation mechanism (how) - is the activity labeled manually, semi-automatically or fully-automatically. All labeling methods have their own benefits, and costs and come with a trade-off between required time and label accuracy. However, not every method is suitable for long-term and in-the-wild recording data. [21], evaluated the annotation performance between 6 different human annotators of a MoCap (Motion Capturing) and IMU HAR Dataset for industrial deployment. They came to the conclusion that annotations were moderately consistent when subjects labeled the data for the first time. However, annotation quality improved after a revision by a domain expert. In the following, we would like to go into more detail on what we consider to be the most important labeling methods for the specific field of activity recognition.
1.1 Annotation Methods in Activity Recognition
Self-Recall methodologies are generally called methods in which study participants have to remember an event in the past. This methodology is used, for instance, in the medical field (e.g. in the diagnosis of injuries [22]), but also frequently in studies in the field of long-term activity recognition. [23] used this method during a study in which participants were asked to label their personal daily data at the end of the day. They noticed that the label quality depends heavily on the participant’s recall and can therefore be very coarse. During a study conducted by [24], every 15 minutes a questionnaire was triggered in which participants needed the answer multiple choice questions about which of 35 predefined activities were recently performed.
App Assisted Labeling: [8] presented in 2014 the so called Prompted labeling. An approach that is already used by commercial smartwatches like the Apple Watch333https://www.apple.com/watch/. In this study user’s were asked to set a label for a time period which has been detected as an activity right after the activity stops. [7] leverages freely available Bluetooth Low Energy (BLE) information broadcasted by other nearby devices and combines this with wearable sensor data in order to detect context and direction changes. The participant is asked to set a new label whenever a change in the signal is detected. [25] published in 2017 the SHL dataset which contains versatile labeled multimodal sensor data that has been labeled using an Android application that asked the user to set a label whenever they detected a position change via GPS. [26] presented a smartphone app that was used in their experimental smart home environment with which study participants were able to either use voice-based labeling, select a label from a list of activities ordered by the corresponding location or scan NFC tags that were installed at locations in the smart house. Similar to [26], [13] developed an open-source mobile app for recording sensor measurements in combination with a self-reported behavioral context (e.g. driving, eating, in class, showering). 60 subjects participated in their study. The study found that most of the participants preferred to fill out their past behavior through a daily journal. Only some people preferred to set a label for an activity that they are about to do. [27] developed a web-based GUI which can either be used on a smartphone, tablet, or a PC to label data recorded in a smart home environment. However, it is important to mention that, According to [8], the process of continually labeling data becomes laborious for participants and can result in a feeling of discomfort.
Unsupervised Labeling is a methodology that uses clustering algorithms to first categorize new samples without deciding yet to which class a sample belongs. [28] presented in 2002 the concept of finding multiple subsets of eigenspaces where, according to [29], each of them corresponds to an individual activity. Huynh uses this knowledge to develop the eigenspace growing algorithm, whereby, growing refers to an increasing set of samples as well as to increasing the so-called effective dimension of a corresponding eigenspace. Based on the reconstruction error (when a new sample is projected to an eigenspace), the algorithm tries to find the best-fitting representation of a sample with minimal redundancy.
[30] recently published a methodology that uses the Pearson Correlation Coefficient to map very specific labels of a variety of datasets to 4 meta labels (inactive, active, walking, and driving) of the ExtraSensory Dataset [13].
Human-in-the-Loop (Labeling) is a collective term for methodologies that integrates human knowledge into their learning or labeling process.
Besides of being applied in HAR research, such techniques are often used in Natural Language Processing (NLP) and according to [31] the NLP community distinguishes between entity extraction [32, 33], entity linking [34], Q&A tasks [35] and reading comprehension tasks [36].
Active Learning is a machine learning strategy that currently receives a lot of attention in the HAR community. Such strategies involves a Human-in-the-Loop for labeling purposes. In the first step the learning algorithm automatically identifies relevant samples of a dataset which are posteriorly queued to be annotated by an expert. Incorporating a human guarantees high quality labels which directly leads to a better performing classifier. Whether a sample is determined to be relevant, and as well the decision to whom it may get presented for annotation purposes are the main focus of research in this field. [37] presents a technique that relies on specific criteria defined by 3 different uncertainty-based selection functions to select samples that will be presented to an expert for labeling and then be propagated throughout the most similar samples. [38] benchmarks the performance of different Active Learning strategies and compared them, with regard to 4 different datasets with a fully-supervised approach. The authors came to the conclusion that Active Learning needs only 8% to 12% of the data to reach similar or even better results than a fully-supervised trained model. These results suggest that presenting pre-selected samples to a human for labeling purposes can reduce the amount of data needed to train a machine learning classifier significantly due to the increased quality of the labels. [39] presented a system which used the Online Active Learning approach published by Scully [40] to bootstrap [41] a machine learning classifier. The publication presented a smartphone app that asked the user right after finishing an activity, which activity has been performed. Afterwards a small subset of the labeled data was used to bootstrap a personalized machine learning classifier.
2 Methodology
Our study is conducted with 11 participants, from which 10 are male and 1 is female. The participants are between 25 and 45 years old. Out of 11 participants, 6 are researchers in the field of signal processing and are used to read and work with sensor data. Participants were selected among acquaintances and colleagues.
2.1 Study Setup
@@
The study was conducted over a period of 2 weeks, during which participants wore an open-source smartwatch on their chosen wrist. Throughout the two-week study period, the participants were instructed to use 4 different labeling methods in parallel, as illustrated in Figure 1. In the first week they were asked to use the in situ button, in situ app, and pure self-recall methods. At the beginning of the 2nd week, we expanded the number of annotation methods with the time-series recall. This annotation method combines the activity diary with a graphical visualization of the participants’ daily data.
The Bangle.js smartwatch has 3 mechanical buttons on the right side of the case. These buttons are programmed to record the number of consecutive button presses per minute. The button-press annotation method captures the total number of button presses along with their corresponding timestamps, enabling the delineation of the beginning and end of an activity within the time-series data. However, this approach does not inherently assign a specific label or description to the identified activity segment. To address this limitation, we employed an inference strategy that leveraged the temporal alignment of the button-press data with the annotations obtained from other methods. By identifying segments with similar timestamps across multiple annotation modalities, we could infer the appropriate label for the button-press annotations.
@@
The time-series recall method can be seen as a combination of an activity diary and a graphical representation of the raw sensor data. For visualization and labeling purposes, we provided the participants with an adapted version of the MAD-GUI. The GUI was published by [9] in 2022 and is a generic open-source Python package. Therefore, it can be integrated into one’s project. Our adaptions to the package are available for download from a GitHub repository444https://github.com/ahoelzemann/mad-gui-adaptions/. It contains changes to the data loader, the definition of available labels, and color settings for displaying the 3D raw data.
Annotation Guidelines: The participants were provided with guidelines that instructed them to document recurring daily activities, encompassing both sports and activities of daily living, that exceeded a duration threshold of 10 minutes. However, the annotation process was deferred until approximately 2 hours after the cessation of the recording session. This temporal offset was implemented to allow for a reasonable time buffer, enabling participants to consolidate their experiences. For instance, if a daily recording concluded at 7 pm, the participant would typically annotate their data around 9 pm, allowing for a 2-hour interim period. Each of these annotation methods represents a layer of annotation that is used for the visual, statistical, and deep learning evaluation. Figure 1 illustrates the overall concept.
Annotation Process: To capture realistic daily data reflecting participants’ natural routines, we granted them complete autonomy in choosing their activity classes. Participants were not restricted to a specific activity protocol; instead, we left the decision of what to label entirely to their judgment. During the study’s first week, participants employed methods - concurrently. In the second week, method was introduced for them to utilize alongside the existing methods. The labels provided by the participants were later interpreted by the researchers and, when necessary, categorized into meta-classes. However, whenever a participant was specific about the activity performed, their label was not summed up into a meta-class. For example, activities such as yoga, badminton, or horse_riding were not combined under the meta-class sport.
2.2 Hardware
Participants wore the commercial open-source smartwatch Bangle.js Version 1 with our open-source firmware555Our smartwatch firmware is made publicly available at: https://github.com/kristofvl/BangleApps/tree/master/apps/activate installed. The device comes with a Nordic 64MHz nRF52832 ARM Cortex-M4 processor with Bluetooth LE, 64kB RAM, 512kB on-chip flash, 4MB external flash, a heart rate monitor, a 3D accelerometer, and a 3D magnetometer. Our firmware only uses the 3D accelerometer and provides the user with the basic functions of a smartwatch, like displaying the time and counting steps. The data is recorded with 25Hz, a sensitivity of ±8g and saved on the devices’ memory with a delta compression algorithm. Therefore, we are able to save up to 8-9 hours (depending on how much of the data could be compressed) of data with the given parameters. The smartwatch stops recording as soon as the memory is full. At the end of the day, the participants need to upload their daily data and program the starting time for the next day using our upload web-tool666Our web-tool is made publicly available at: https://ubi29.informatik.uni-siegen.de/upload/.
@@ 3 Statistical Analysis
The labels were statistically analyzed based on their consistency using the Cohen score as well as the number of missing annotations across all methods. The Cohen score describes the agreement between two annotation methods, which is defined as follows , see [42] and [21]. Where is the observed agreement ratio and is the expected agreement if both annotators assign labels randomly. The score shows how uniform two different annotators labeled the same data. For calculation purposes, an implementation provided by Scikit-Learn [43], was used. Furthermore, missing annotations across methods are measured as the percentage of missing or incomplete annotations. The annotations of all methods were first compared with each other and matched based on the given time indications. Annotations that could not be assigned or were missing were marked accordingly and are the base for calculating this indicator, see Section 5.2 for more information.
We used a similar representation as [44] to visualize the matches among labeling methods. In this study, the authors compared genome annotations labeled by different annotators with regard to their error scores between different annotators.
@@ 4 Effects on Deep Learning Performance
The implementation of [45] incorporates BatchNormalization layers after each Convolutional layer, as well as MaxPooling for the transition between the final convolutional block and the LSTM layer, and a Dropout layer before classification. Each Convolutional layer employs a ReLU activation function. The inclusion of the BatchNormalization layers serves to accelerate training and mitigate the detrimental effects of internal covariate shift, as discussed further in [49].
Preprocessing: To prepare the data for neural network training, we perform two preprocessing steps. First, we address minor inconsistencies in the device’s sampling rate. The original data was collected at a rate of 12.5 Hz. However, for optimal performance with neural networks, a consistent and regular sampling rate is preferred. To achieve this, we upsample the data by a factor of two, resulting in a constant frequency of 25 Hz. This upsampling process essentially inserts additional data points between the existing ones, effectively increasing the resolution of the signal. The second preprocessing step involves rescaling the accelerometer data to a range between -1 and 1.
Leave-One-Day-Out Cross-Validation: Figure 3 illustrates the train and test setting for the deep learning model. Instead of following the traditional Leave-One-Subject-Out strategy, we adapted it to our needs by using one day of the week for testing and training on the remaining days for each study participant and week.
@@
This approach was necessitated by the unique characteristics of our dataset. It consists of a predominant void class and a small number of samples per activity class and participant. To mitigate the issue of an disproportionately large void class, we trained our model with balanced class weights. By not limiting the participants in their choice of daily activities and not specifying predefined activity labels, we ended up with very unique sets of activities for each study participant. Given these circumstances, it is unrealistic to expect a model capable of generalizing across participants and days. The Leave-One-Day-Out strategy aims to maintain the consistency of class labels within each day’s data, providing a more cohesive and reliable dataset for training and evaluation purposes. This strategy also mitigates the potential impact of participant-specific biases or variations in class labeling, leading to a more robust and accurate model. Furthermore, due to the in-the-wild recording setup, the intra-class differences [50] for comparatively simple activities, such as walking or running can be significant. Consequently, the impact of different labeling methods is expected to be more pronounced and visible in a personalized model compared to a generalized model.
Post-Processing & Classification: In the classification task, we initially segmented the data into fixed-length sliding windows of 2 seconds (50 samples). However, our objective extended beyond instantaneous classifications; we aimed to identify longer periods of recurring activities. To achieve this, we employed a post-processing technique involving a jumping window approach with a duration of 5 minutes. Within each 5-minute window, a majority vote was applied to the individual 2-second window predictions. The activity class with the highest number of occurrences within the 5-minute window was then assigned as the predominant activity for the entire window. This approach enabled us to capture sustained patterns of activities over extended periods, aligning with our goal of analyzing longer-term behavioral trends.
@@ 5 Results
Our participants were asked to annotate daily activities lasting more than 10 minutes. We did not limit them to a predefined set of classes; they independently decided on labels for their activities. After normalizing the labels (e.g., changing ”going for a walk” to ”walking”), the participants assigned 26 different labels: laying, sitting, walking, running, cycling, bus_driving, car_driving, cleaning, vacuum_cleaning, laundry, cooking, eating, shopping, showering, yoga, sport, playing_games, desk_work, guitar_playing, gardening, table_tennis, badminton, horse_riding, cleaning, reading, weightlifting, manual_work, dish_washing. Any unlabeled samples were classified as void. However, after excluding infrequent or non-standalone classes (e.g., shopping likely combines walking, standing, and sitting), we reduced the dataset to 23 labels (22 activities plus void): laying, sitting, walking, running, cycling, bus_driving, car_driving, vacuum_cleaning, laundry, cooking, eating, shopping, showering, yoga, sport, playing_games, desk_work, guitar_playing, gardening, table_tennis, badminton, horse_riding. Nevertheless, the graphical representation of the distribution and the table in Section 5.1 include the full scope of classes.
5.1 Class Distribution
The class distribution reflects a broad range of activity classes that represent the daily lives of our participants.
These classes remain primarily participant-specific due to the absence of a predefined annotation protocol, which allows participants the freedom to label activities according to their own interpretations. Consequently, we decided against employing a Leave-One-Subject-Out evaluation method, as it might introduce inconsistencies in the dataset due to the varying class labels assigned by different participants. The walking class is the most consistently annotated class across participants and annotation methods, although it may not represent the maximum amount of labeled data points. Notably, the void class, which is not visible in Figure 4, accounts for a substantial portion of the labeled data, ranging from 80% to 96%, depending on the annotation method employed.
Table 1 shows that daily activities that do not require extensive planning and are inherent to most people’s everyday lives tend to be the most consistently annotated classes. Among these are activities such as walking, cycling, car_driving, or cooking. On the other hand, activities like badminton, weightlifting, manual_work, and others are highly subject-dependent and occur only sporadically. The observation that the desk_work class exhibits the highest frequency is valid solely when employing the diary- or the GUI-methodology for data collection, suggesting a potential limitation or bias associated with this particular approach. The classes pertaining to physical activities such as running, bus_driving, yoga, badminton, weightlifting, and sport exhibit a distinct pattern of clustering within specific weeks, indicating a temporal dependency. This observation highlights the inherent bias introduced by the real-world recording environment in which the dataset was collected, potentially limiting the generalizability of the model to broader contexts. Furthermore, it is crucial to note that the size of the void class for the recall methods and is up to 16% smaller than the in situ methods and . While this disparity in class representation highlights the inherent complexities and challenges associated with the data collection and annotation procedures, it simultaneously presents an opportunity to address real-world imbalances and biases. By critically examining and accounting for these factors, the resulting models can potentially enhance their generalizability and applicability across diverse scenarios, ultimately contributing to a deeper understanding of the underlying phenomena under investigation.
Week 1 Week 2 Instances button app diary button app diary GUI Total laying 0 0 135,000 0 0 0 0 1 sitting 0 0 292,500 0 0 607,500 264,041 13 walking 1,409,177 1,380,112 2,462,923 1,446,000 1,168,525 2,160,712 2,228,337 245 running 16,500 74,600 84,000 0 0 0 0 4 cycling 573,000 574,100 647,880 616,500 330,000 610,500 784,224 92 bus_driving 0 67,500 15,000 0 0 0 0 2 car_driving 708,000 659,950 1,683,176 340,500 155,450 1,025,985 954,882 111 vacuum_cleaning 49,500 66,000 67,500 0 42,000 60,000 47,955 10 laundry 0 27,000 30,000 33,000 76,500 112,500 106,316 19 cooking 229,500 45,632 269,132 214,500 96,000 382,500 419,848 44 eating 0 0 5,753 0 0 75,000 92,850 5 shopping 0 57,000 60,000 0 0 45,000 0 3 showering 112,500 0 225,000 34,500 0 251,777 165,721 20 yoga 0 0 0 0 0 22,500 30,680 2 sport 39,000 46,088 45,000 0 0 0 0 3 playing_games 0 0 648,926 0 0 45,000 100,771 6 desk_work 36,000 0 3,646,985 0 0 4,537,832 1,137,214 21 guitar_playing 49,500 0 172,500 111,000 67,060 217,500 234,355 24 gardening 48,000 53,125 43,718 69,000 69,650 375,000 421,794 16 table_tennis 0 17,875 0 190,500 108,725 105,000 171,177 29 badminton 0 0 0 0 234,000 210,000 252,725 6 horse_riding 0 66,000 199,500 0 519,000 495,000 502,742 18 cleaning 70,500 100,500 240,000 48,000 40,500 502,500 215,034 25 reading 0 0 0 0 0 45,000 0 1 weightlifting 0 0 0 84,000 0 60,000 0 2 manual_work 0 0 0 0 0 69,425 157,587 3 dish_washing 0 0 0 10,500 21,000 52,500 66,246 9 void 61,581,367 61,685,562 53,948,051 60,240,422 60,510,012 51,369,691 55,083,923 539